Evaluate the functions for the values of x given as 1, 2, 4, 8, and 16. Describe the differences in the rate at which each function changes with increasing values of x. And then rank them fastest to slowest using the problem numbers.
f(x) = 2x – 5
f(x) = x2 - 4x + 3
f(x) = x3 + 4x2 - 2x - 3
f(x) = 7x
f(x) = log x

Found 2 solutions by solver91311, erie_acebedo:
Answer by solver91311(24713)



Just substitute the given numbers for x and do the arithmetic. I hope you didn't expect someone to do these 25 calculations for you.

John


Answer by erie_acebedo(2)

f(x)=2x-5
y=2x-5
x=2y-5
2y/2=x-5/2
y=x-5/2